OpenStudy (anonymous):

Solve C = 3(2f – 5h) for f in terms of C and h. You must show all work.

OpenStudy (mathstudent55):

C = 3(2f - 5h) To solve for f means to isolate f. First we distribute 3 on the right side: C = 3(2f) - 3(5h) C = 6f - 15h Now that we have f in a sinlge term, let's get rid of -15h by adding 15h to both sides C + 15h = 6f Switch sides: 6f = C + 15h Since f is being multiplied by 6, we divide both sides by 6: f = (C + 15f)/6